Solve the system y = -x + 7 and y = -0.5(x - 3)^2 + 8

Answer by jorel1380(3719)

If both equations equal y, then
-x+7=-0.5(x-3)²+8
-0.5(x-3)²+x+1=0
(x-3)²-2x-2=0☺
x²-8x+7=0
(x-7)(x-1)=0
x=7 or 1
So the sets are (7,0) and (1,6). ☺☺☺☺